OpenStudy (anonymous):

6z^3 over 7t^3 multiplied by 49t^7 over 36z? Anyone??

OpenStudy (anonymous):

((6z^2)/(7t^3))((49t^7)/(36z)) is that the question

OpenStudy (anonymous):

start with \[\frac{6z^3\times 49t^7}{7t^3\times 36z}\] then cancel your brains out

OpenStudy (anonymous):

no, by "over" I mean like a fraction.... 6zx^2/7t^3

OpenStudy (anonymous):

get \[\frac{7z^2t^4}{6}\]

OpenStudy (anonymous):

OMG TY tyt yty...100% me luv u long time ;D